# Towards Faithful and Plausible Natural Language Explanations for Image Classification: A Pipeline Approach

### v1

@@ Abstract
Existing explanation methods for image classification struggle to provide faithful and plausible explanations. This paper addresses this issue by proposing a post-hoc natural language explanation method that can be applied to any CNN-based classifier without altering its training process or affecting predictive performance. By analysing influential neurons and the corresponding activation maps, the method generates a faithful description of the classifier’s decision process in the form of a structured meaning representation, which is then converted into text by a language model. Through this pipeline approach, the generated explanations are grounded in the neural network architecture, providing accurate insight into the classification process while remaining accessible to non-experts. Experimental results show that the NLEs constructed by our method are significantly more plausible and faithful. In particular, user interventions in the neural network structure (masking of neurons) are three times more effective than the baselines.
@@
Adam Wojciechowski1and Mateusz Lango1,2and Ondřej Dušek2 1 Poznan University of Technology, Faculty of Computing and Telecommunications, Poznan, Poland 2 Charles University, Faculty of Mathematics and Physics, Prague, Czech Republic {lango,odusek}@ufal.mff.cuni.cz
@@ 1 Introduction
Therefore, several types of explanation methods have been proposed, including various forms of saliency maps Selvaraju et al. (2017a), feature importances Ribeiro et al. (2016a), concept-based explanations Chen et al. (2019), counterfactual explanations Vermeire et al. (2022), etc. A particularly interesting form of explaining predictions is offered by natural language explanations (NLE) techniques (Camburu et al., 2018; Wu and Mooney, 2019). Such explanations are not only understandable by non-expert users, but can also be used to support conversations with the user in a dialogue system Raczyński et al. (2023).
@@
In the case of NLE, obtaining high plausibility is straightforward, as textual explanations are usually human-friendly Gurrapu et al. (2023), but achieving high fidelity is challenging. In the context of image classification, image captioning methods offer plausible but unfaithful NLEs Xu et al. (2015); Kamakshi and Krishnan (2023). Some methods try to improve faithfulness by conditioning generation on both the predicted class and image features Hendricks et al. (2016); Kim et al. (2018); Marasović et al. (2020); Sammani et al. (2022), but the faithfulness provided is still limited as the model is not aware of the classifier’s decision process. Other methods train image classifiers to jointly predict the class and visual rationales, and generate explanations based on them Wickramanayake et al. (2021); Kayser et al. (2022). However, the rationales are predicted independently of the class and do not participate in the classifier’s decision process. Most importantly, such methods change the training procedure and the architecture of the classifier, often affecting classification performance.
In this paper, we propose a post-hoc natural language explanation method for image classification that can be used with any standard convolutional neural network (CNN) classifier. To illustrate the classification decision process, the method analyses which neurons of the CNN were most influential in reaching a given decision, and which regions of the image caused them to activate. For each influential neuron, a neuron annotation method computes visual pattern exemplars and produces a short (one phrase) description of a pattern that the neuron detects. The information gathered in this process serves to create a simple, human-readable meaning representation, which is converted to text by a large language model (LLM).
Our simple two-step pipeline method (see Fig. 1) produces explanations that are directly grounded in the classifier’s network architecture, but without interfering with its training process or affecting its predictive performance. It also does not need any gold-standard explanations. The provided NLEs reflect the underlying classification process by specifying the most influential neurons computed by a well-established explainable AI methods. At the same time, the final text is fluent and easy to understand. This results in NLEs that are both plausible and highly faithful. Our experimental code is publicly available. [^1] http://tinyurl.com/3da45hnm
@@ 2.1 Meaning representation construction
To choose the most relevant neurons, we apply well-established Layer-wise Relevance Propagation (LRP) method Bach et al. (2015). LRP performs a backward pass through the classifier network to establish the influence of each neuron to the final prediction (see App. I for formulas). We select $k$ neurons with the highest LRP scores (with $k$ being a parameter controlling the brevity-detail tradeoff).
@@
The neuron’s raw activation map is divided into a 3x3 grid with manually assigned labels such as ‘top-left corner’, ‘top’, ‘top-right corner’, etc. We then select all grid cells where the neuron’s activation exceeded half of its maximum value. We apply several substitution rules (see App. C) to make the list of cells shorter and more human-readable.
@@ 2.2 Explanation generation
The second step of our method is converting the human-readable and faithful MR created above into a user-friendly text. As we do not have any gold-standard explanation texts, the task is performed by prompting a large language model (LLM).
We instructed the model to (1) produce fluent text, (2) summarise the content of the MR (e.g. if two neurons detect similar patterns, they can be combined in the text), (3) prioritise readability, (4) come up with its own formulation of spatial positions to improve fluency. We also provide one handcrafted MR-to-text conversion example. The prompt is shown in App. D. LLMs could in theory hallucinate and thus reduce the explanations’ faithfulness. However, in Section 3 we show experimentally that current LLMs are reliable enough to produce useful explanations.
@@ 3.1 Experimental setup
All experiments were performed on the ImageNet dataset. The classifier was trained on the train set and our explanation method was run to explain predictions made on the validation data. [^2] Annotated ImageNet test set is not publicly available.
@@
We experiment with explaining the predictions of the smallest CNN classifier from the popular ResNet family: ResNet18 He et al. (2015).333 It reaches only a 41.4% accuracy on the validation set, but high classification performance is not the goal of our study. We fill our MRs with $k=10$ top neurons indicated by LRP from the Captum library Kokhlikyan et al. (2020) and annotate them using MILAN’s original implementation. As the LLM for MR-to-text conversion, we use GPT-4 Achiam et al. (2023).
@@ 3.2 Are the output explanations plausible?
To assess the plausibility of generated explanations, we conducted a small-scale manual annotation experiment. We recruited ten annotators: five non-experts hired on the Prolific platform and five experts with at least one published paper on explainable AI. Each annotator was presented with 30 image-explanation pairs (300 in total) and asked to rate them on a scale of 1-5 whether the explanation was (1) fluent, (2) easy to understand (comprehensible), (3) convincing, and (4) explanatory for the underlying decision process. [^4] Note that the question on understanding the decision process does not measure faithfulness, but the user’s subjective opinion on whether they understand how the model works. The overall quality of the explanations was also rated (see App. E).
The results are presented in Tab. 1. Our method obtains the highest overall quality according to both experts and non-experts. It also produces the most plausible explanations (most convincing and explanatory). Since the baselines produce much shorter explanations, it is not surprising that our longer explanations are a bit more difficult to understand. Interestingly, experts generally give higher ratings than non-experts. This is true for all methods and all factors except the explanatory ability. For this factor, experts rate the baselines lower than non-experts, but they consistently rate the explanations provided by our pipeline higher. The improvements of our method over baselines are statistically significant on both plausibility measures and overall quality. For fluency, our method is indistinguishable from SAT (see details in App. J).
@@ 3.3 Are the output explanations faithful?
We asked annotators to cover with white rectangles parts of images that contained the decision rationale indicated in the NLE, 50% area at most (see App. H for details). We re-classified covered images and measured changes in prediction and the average decrease in the probability of the originally predicted class. We also performed an opposite experiment, with the annotators highlighting only parts of image mentioned in the explanation and covering the rest.
@@
We also re-ran our NLE pipeline with parts of the input image covered. This led to significant changes: on average 78% (median 90%) of the neurons indicated in MRs were different.
@@
To show the NLEs’ ability to reflect classifier decisions, we asked the annotators to read the NLEs and select up to five most influential neurons from a MILAN-annotated list. The classifier was then re-run with the selected neurons masked. The results in Table 2 reveal that masking neurons suggested by our NLEs led to a five times higher decrease in the predicted class probability and over three times higher class flip rate than baselines. More details are in App. B.
@@
Following Wiegreffe et al. (2021), we measure explanation robustness against adding random noise to the input image (intra-set stability, see App. G for details) by comparing BLEU Papineni et al. (2002), METEOR Lavie and Agarwal (2007) as well as class flip frequency and class probability change against the original predictions. We also check for outputs’ diversity (inter-set stability) using BLEU and METEOR overlap against explanations for other classes. The results in Table 3 show high diversity and high noise sensitivity: In parallel to increased classification changes with added noise, explanation BLEU/METEOR gradually drop, but they are lowest when comparing between different classes.
@@
The human evaluation of our approach’s MR-to-text reliability was similar to plausibility evaluation, but limited to non-expert Prolific annotators. We asked five yes-no questions on information in the text not grounded in the MR (i.e., hallucinations), omission of MR information, fluency, spatial information fidelity, and overall correctness (see App. F).
@@ Appendix A Examples of explanations
Several examples of explanations provided by the methods under study are given in Tab. 4. More examples of NLE produced with the proposed pipeline approach are given in Tab. 5.
@@
Appendix B Detailed results of masking influential neurons experiment
The detailed results of the experiment with masking influential neurons are presented in Fig. 2. For our method, the results indicate that masking neurons in the order indicated by the annotators leads to an increasing change in the classifier’s prediction and a larger decrease in the predicted class probability value. In contrast, the annotators gain limited insight into how the neural network made a decision from the explanations provided by other methods: masking the indicated neurons leads to limited change, and increasing the number of masked neurons leads to almost no improvement. We attribute the effectiveness of masking the first indicated neuron to examples of classes that are highly correlated with a pattern detected by a particular neuron.
(a) The percentage of changed predictions after masking selected neurons
(b) The decrease of the probability of the predicted class after masking selected neurons
Figure 2: The results of neuron masking experiments for varying number of masked neurons.
Appendix C Simplification rules for spacial positions
@@
Appendix D Prompt for MR-to-text conversion
To convert meaning representations into text, the following prompt was applied to the language model: {spverbatim} You are given a problem of creating textual explanation of an image classification performed by neural network. You will be given a Python object representing network output in the form of ‘’image class’, [(’detected object’, ’position’), …]‘. I want you to convert this object into a textual explanation. You should: 1. Create a grammatically correct sentence which will explain the model’s decision. 2. Decide which detected objects do not fit with image class and do not include them in the explanation. For instance, ’dentist’ class and ’animal heads’ objects are completely unrelated. However, the descriptions that aren’t directly related to image class, but can be indirectly correlated, especially in terms of shape, color, or texture resemblance should be included (like ’fountain’ and ’sea’ because of the water they have in common or ’brick wall’ and ’grid’ because the texture is similar). Never mention that you chose neuron descriptions and do not talk about the neuron descriptions that were discarded. 3. Prioritize the readability of the explanation. Include only essential detected objects and aggregate information to shorten the explanation. 4. Aggregate positions if possible, for example [’bottom-left corner’, ’bottom’, ’bottom-right corner’] should be aggregated into ’bottom’. If the positions list is too long or too ambiguous do not include them in the explanation.
@@ Appendix D Prompt for MR-to-text conversion
Appendix E Human evaluation of explanation plausibility
The annotators are presented with an image, model’s prediction and a natural language explanation. Each question is answered on a scale from 1 (low) to 5 (high). The following binary questions are asked:
@@ Appendix E Human evaluation of explanation plausibility
Appendix F Human evaluation of MR-to-text transformation
@@
  • If any spatial compression occurred between explained neurons, the said compression is correct?
  • Overall, do you find this meaning representation to text transformation acceptable i.e. sufficiently good for explanation purposes?
@@
Image Class Explanation
[Uncaptioned image] Target: "lakeside", Prediction: "lakeside" The model classified this image as ’lakeside’ because it detected water in the lower half and the bottom-right corner of the image. It also noticed volcanoes and mountains in the upper half, and walkways at the bottom-left corner and bottom. Additionally, the model found items with straight features at the left, bottom-left corner, and bottom of the image.
[Uncaptioned image] Target: "coral reef", Prediction: "coral reef" The model classified this image as a "coral reef" due to the detection of aquatic life throughout the image, particularly at the bottom, left, and right. The presence of water covering the entire image and nature-related elements further supported this classification. The model also noticed items that are connected in the lower half of the image and similar color patterns, which are common in coral reef environments.
[Uncaptioned image] Target: "wall clock", Prediction: "wall clock" The model classified this image as a "wall clock" because it detected clocks and other gauges, circular, round objects, and black and white objects at the bottom-right corner, center, and bottom of the image. It also noticed rounded edges in the right half of the image and the indent in an hourglass shape at the bottom-right corner, center, right, and bottom.
Table 5: Examples of explanations provided by the method
Appendix G Details on stability experiments
Let us assume that an image is a matrix $X$, such that $X\in[0,1]$. We model input perturbations by adding random noise to the images from the standard normal distribution. Since an interval of possible pixel values is $[0,1]$, to account for the unboundedness of a standard normal distribution we use a clipping operation defined as follows:

### v2

@@ Abstract
Existing explanation methods for image classification struggle to provide faithful and plausible explanations. This paper addresses this issue by proposing a post-hoc natural language explanation method that can be applied to any CNN-based classifier without altering its training process or affecting predictive performance. By analysing influential neurons and the corresponding activation maps, the method generates a faithful description of the classifier’s decision process in the form of a structured meaning representation, which is then converted into text by a language model. Through this pipeline approach, the generated explanations are grounded in the neural network architecture, providing accurate insight into the classification process while remaining accessible to non-experts. Experimental results show that the NLEs constructed by our method are significantly more plausible and faithful than baselines. In particular, user interventions in the neural network structure (masking of neurons) are three times more effective.
@@
Adam Wojciechowski1,3and Mateusz Lango1,2and Ondřej Dušek2 1Poznan University of Technology, Faculty of Computing and Telecommunications, Poland 2Charles University, Faculty of Mathematics and Physics, Prague, Czechia 3Samsung AI Center Warsaw, Poland a.wojciecho4@samsung.com, {lango,odusek}@ufal.mff.cuni.cz
@@ 1 Introduction
Therefore, several types of explanation methods have been proposed, including various forms of saliency maps Selvaraju et al. (2017), feature importances Ribeiro et al. (2016), concept-based explanations Chen et al. (2019), counterfactual explanations Vermeire et al. (2022), etc. A particularly interesting form of explaining predictions is offered by natural language explanation (NLE) techniques (Camburu et al., 2018; Wu and Mooney, 2019). Such explanations are not only understandable by non-expert users, but can also be used to support conversations with the user in a dialogue system Raczyński et al. (2023).
@@
In the case of NLE, obtaining high plausibility is straightforward, as textual explanations are usually human-friendly Gurrapu et al. (2023), but achieving faithfulness is challenging. In the context of image classification, image captioning methods offer plausible but unfaithful NLEs Xu et al. (2015); Kamakshi and Krishnan (2023). Some methods try to improve faithfulness by conditioning generation on both the predicted class and image features Hendricks et al. (2016); Kim et al. (2018); Marasović et al. (2020); Sammani et al. (2022), but the faithfulness provided is still limited as the model is not aware of the classifier’s decision process. Other methods train image classifiers to jointly predict the class and visual rationales, and generate explanations based on them Wickramanayake et al. (2021); Kayser et al. (2022). However, the rationales are predicted independently of the class and do not participate in the classifier’s decision process. Most importantly, such methods change the training procedure and the architecture of the classifier, often affecting classification performance.
In this paper, we propose a post-hoc natural language explanation method for image classification that can be used with any standard convolutional neural network (CNN) classifier. To illustrate the classification decision process, the method analyses which neurons of the CNN were most influential in reaching a given decision, and which regions of the image caused them to activate. For each influential neuron, a neuron annotation method computes visual pattern exemplars and produces a short (one phrase) description of a pattern that the neuron detects. The information gathered in this process serves to create a simple meaning representation, which is converted to natural-language text by a large language model (LLM).
Our simple pipeline method (see Fig. 1 and Sect. 2) produces explanations that are directly grounded in the classifier’s network architecture, but without interfering with its training process or affecting its predictive performance. It also does not need any gold-standard training explanations. The provided NLEs reflect the process underlying classification by specifying the most influential neurons computed by well-established explainable AI methods. At the same time, the final text is fluent and easy to understand. This results in NLEs that are both plausible and highly faithful – significantly more so than baselines, as demonstrated in Sect. 3. Our experimental code is publicly available. [^1] https://github.com/wojciechowskiofficial/FLEX
@@ 2.1 Meaning representation construction
To pick the most relevant neurons, we apply the well-established Layer-wise Relevance Propagation (LRP) method Bach et al. (2015). LRP performs a backward pass through the classifier network to establish the influence of each neuron to the final prediction (see App. H for formulas). We select $k$ neurons with the highest LRP scores (with $k$ being a parameter controlling the brevity-detail tradeoff).
@@
The neuron’s raw activation map is divided into a 3$\times$3 grid with manually assigned labels such as ‘top-left corner’, ‘top’, ‘top-right corner’, etc. We then select all grid cells where the neuron’s activation exceeded half of its maximum value. We apply several substitution rules (see App. B) to make the list of cells shorter and more human-readable.
@@ 2.2 Explanation generation
The second step of our method is converting the faithful MR created above into a user-friendly text. As we do not have any gold-standard explanation texts, the task is performed by prompting a large language model (LLM).
We instructed the model to (1) produce fluent text, (2) summarise the content of the MR (e.g. if two neurons detect similar patterns, they can be combined in the text), (3) prioritise readability, (4) come up with its own formulation of spatial positions to improve fluency. We also provide one handcrafted MR-to-text conversion example. The prompt is shown in App. C. LLMs could in theory hallucinate and thus reduce the explanations’ faithfulness. However, in Section 3 we show experimentally that current LLMs are reliable enough to produce useful explanations.
@@ 3.1 Experimental setup
All experiments were performed on the ImageNet dataset. The classifier was trained on the training set and our explanation method was run to explain predictions made on the validation data. [^2] Annotated ImageNet test set is not publicly available.
@@
We experiment with explaining the predictions of the smallest CNN classifier from the popular ResNet family: ResNet18 He et al. (2015).333 It reaches only a 41.4% accuracy on the validation set, but high classification performance is not the goal of our study. We fill our MRs with $k=10$ top neurons indicated by LRP from the Captum library Kokhlikyan et al. (2020) and annotate them using MILAN’s original implementation. As the LLM for the MR-to-text conversion, we employ GPT-4 (gpt-4-0613; Achiam et al., 2023).
@@ 3.2 Are the output explanations plausible?
To assess the plausibility of generated explanations, we conducted a small-scale manual annotation experiment. We recruited ten annotators: five non-experts hired on the Prolific platform and five experts with at least one published paper on explainable AI. Each annotator was presented with 30 image-explanation pairs (300 in total) and asked to rate on a scale of 1-5 whether the explanations were (1) fluent, (2) easy to understand (comprehensible), (3) convincing, and (4) insightful for the underlying decision process. [^4] Note that the question on understanding the decision process does not measure faithfulness, but the user’s subjective opinion on whether they understand how the model works. The overall quality of the explanations was also rated (see App. D).
The results are presented in Tab. 1 and examples of generated NLEs can be found in Tab. 3 (see App. A for more). Our method obtains the highest overall quality according to both experts and non-experts. It also produces the most plausible explanations (most convincing and insightful). Since the baselines produce much shorter explanations, it is not surprising that our longer explanations are a bit more difficult to understand. Interestingly, experts generally give higher ratings than non-experts for all methods and all factors except for providing insight into the decision process. Here, experts rate the baselines lower than non-experts, but they consistently rate the explanations provided by our pipeline higher. The improvements of our method over baselines are statistically significant on both plausibility measures and overall quality. For fluency, our method is indistinguishable from SAT (see details in App. I).
@@ 3.3 Are the output explanations faithful?
We asked annotators to cover with white rectangles parts of images that contained the decision rationale indicated in the NLE, 50% area at most (see App. G for details). We re-classified covered images and measured changes in prediction and the average decrease in the probability of the originally predicted class. We also performed an opposite experiment, with the annotators highlighting only parts of image mentioned in the explanation and covering the rest.
@@
We also re-ran our NLE pipeline with parts of the input image covered. This led to significant changes: on average, 78% (median 90%) of the neurons indicated in MRs were different.
@@
To show the NLEs’ ability to reflect classifier decisions, we asked the annotators to read the NLEs and select up to five most influential neurons from a MILAN-annotated list. The classifier was then re-run with the selected neurons masked. The results in Table 2 reveal that masking neurons suggested by our NLEs led to a five times higher decrease in the predicted class probability and over three times higher class flip rate than baselines.
More detailed results are presented in Fig. 3. Masking neurons in the order indicated by the annotators using our method leads to an increasing change in the classifier’s prediction and a gradual decrease in the predicted class probability value. In contrast, masking the neurons indicated using other methods leads to a small decrease in class probability for one masked neuron and almost no further decrease for more masked neurons. As we attribute the effect of the first masked neuron to examples of classes that are highly related to a singular pattern (e.g., a neuron annotated “water” for the class “sea”), this indicates that annotators gain very little insight into how the neural network made a decision from the explanations provided by the baselines.
Figure 2: Results of neuron masking experiments when varying the number of masked neurons. Left: percentage of changed predictions after masking selected neurons; right: decrease of the probability of the predicted class after masking selected neurons.
Image Class Explanation
Refer to caption Target: "lakeside"
Prediction: "lakeside"
The model classified this image as ’lakeside’ because it detected water in the lower half and the bottom-right corner of the image. It also noticed volcanoes and mountains in the upper half, and walkways at the bottom-left corner and bottom. Additionally, the model found items with straight features at the left, bottom-left corner, and bottom of the image.
Refer to caption Target: "coral reef"
Prediction: "coral reef"
The model classified this image as a "coral reef" due to the detection of aquatic life throughout the image, particularly at the bottom, left, and right. The presence of water covering the entire image and nature-related elements further supported this classification. The model also noticed items that are connected in the lower half of the image and similar color patterns, which are common in coral reef environments.
Refer to caption Target: "wall clock"
Prediction: "wall clock"
The model classified this image as a "wall clock" because it detected clocks and other gauges, circular, round objects, and black and white objects at the bottom-right corner, center, and bottom of the image. It also noticed rounded edges in the right half of the image and the indent in an hourglass shape at the bottom-right corner, center, right, and bottom.
Figure 3: Examples of explanations provided by our method.
@@
Following Wiegreffe et al. (2021), we measure explanation robustness against adding random noise to the input image (intra-set stability, see App. F for details) by comparing BLEU Papineni et al. (2002), METEOR Lavie and Agarwal (2007) as well as class flip frequency and class probability change against the original predictions. We also check for outputs’ diversity (inter-set stability) using BLEU and METEOR overlap against explanations for other classes. The results in Table 3 show that the explanations are both distinct for different classes and highly sensitive to noise: As we add more noise resulting in increased classification changes, the explanation BLEU and METEOR gradually drop, but they are at their lowest when comparing between different classes.
@@
The human evaluation of our approach’s MR-to-text reliability was similar to plausibility evaluation, but limited to non-expert Prolific annotators. We asked five yes-no questions on information in the text not grounded in the MR (i.e., hallucinations), omission of MR information, fluency, spatial information fidelity, and overall correctness (see App. E).
@@
Additionally, we repeated this experiment for explanations generated by an open LLM (Llama 3 70B) instead of GPT-4. The results presented in App. E show that open LLM generated NLEs with a significantly higher number of hallucinations and omissions, but this did not affect the overall quality score given by the annotators.
@@ Appendix A Examples of explanations
Several examples of explanations provided by the methods under study are given in Tab. 4.
@@
Appendix B Simplification rules for spacial positions
@@
Appendix C Prompt for MR-to-text conversion
To convert meaning representations into text, the following prompt was applied to the language model:
{spverbatim}
You are given a problem of creating textual explanation of an image classification performed by neural network. You will be given a Python object representing network output in the form of ‘’image class’, [(’detected object’, ’position’), …]‘. I want you to convert this object into a textual explanation. You should: 1. Create a grammatically correct sentence which will explain the model’s decision. 2. Decide which detected objects do not fit with image class and do not include them in the explanation. For instance, ’dentist’ class and ’animal heads’ objects are completely unrelated. However, the descriptions that aren’t directly related to image class, but can be indirectly correlated, especially in terms of shape, color, or texture resemblance should be included (like ’fountain’ and ’sea’ because of the water they have in common or ’brick wall’ and ’grid’ because the texture is similar). Never mention that you chose neuron descriptions and do not talk about the neuron descriptions that were discarded. 3. Prioritize the readability of the explanation. Include only essential detected objects and aggregate information to shorten the explanation. 4. Aggregate positions if possible, for example [’bottom-left corner’, ’bottom’, ’bottom-right corner’] should be aggregated into ’bottom’. If the positions list is too long or too ambiguous do not include them in the explanation.
@@ Appendix C Prompt for MR-to-text conversion
Appendix D Human evaluation of explanation plausibility
The annotators are presented with an image, model’s prediction and a natural language explanation. Each question is answered on a scale from 1 (low) to 5 (high). The following questions are asked:
@@ Appendix D Human evaluation of explanation plausibility
Appendix E Human evaluation of MR-to-text transformation
@@
  • If any spatial compression occurred between explained neurons, is the said compression correct?
  • Overall, do you find this meaning representation to text transformation acceptable, i.e. sufficiently good for explanation purposes?
@@
The experiment was carried out on explanations generated with two LLMs: GPT-4 (used in all other experiments) and an open-weight alternative, namely Llama 3 70B from ollama library [^5] https://ollama.com/library/llama3. Both sets of explanations were generated using the same prompt and for the same MRs.
The results are presented in Tab. 5. Both GPT-4 and Llama appear to have a similar ability to perform spatial compressions, but for the other factors examined, the NLEs generated by Llama fall short of those generated by GPT-4. Llama’s explanations are almost 6 times more likely to contain hallucinations. They also have more omissions and lower fluency. Nevertheless, the overall correctness of the NLE’s generated by both LLMs is the same and not overly high. We think that this result is influenced by the fact that some annotators penalise the correctness of NLE’s too much due to omitted information from the meaning representation. Note that the MRs are quite long and the generated NLEs are supposed to summarise them and shorten them to improve readability, thus omitting some information.
Question GPT-4 Llama 3 70B
Hallucinations $\downarrow$ 0.08 0.46
Omissions $\downarrow$ 0.44 0.64
Fluency $\uparrow$ 0.96 0.86
Spacial compression $\uparrow$ 0.82 0.82
Overall correctness $\uparrow$ 0.58 0.58
Table 5: Human evaluation of MR to text conversion with GPT-4 and Llama 3 70B as backbone LLMs. The percentage of "yes" answers is reported.
Appendix F Details on stability experiments
Let us assume that an image is a matrix $X$, such that $X\in[0,1]$. We model input perturbations by adding random noise to the images, sampled from the standard normal distribution. Since an interval of possible pixel values is $[0,1]$, to account for the unboundedness of a standard normal distribution we use a clipping operation defined as follows: